# An Efficient Dual-Line Decoder Network with Multi-Scale Convolutional Attention for Multi-organ Segmentation

### v1

@@ 4.6 Result on SegThor Dataset
The comparison in Table 3 evaluates segmentation methods for four anatomical structures—Esophagus, Heart, Trachea, and Aorta—using metrics like Dice similarity, ASD, and IoU. AttUNet performs well in ASD for the Esophagus but falls behind in Dice and IoU, which indicates challenges with boundary segmentation accuracy. Similarly, R2UNet struggles across all metrics, especially for the Esophagus and Heart, reflecting its limitations in capturing complex anatomical details. Advanced methods like UNet++ and UDBANet show notable improvements, with UNet++ excelling in Heart and Trachea segmentation (Dice of 0.95 and 0.89) and UDBANet performing strongly for the Heart and Aorta, though its precision for the Esophagus is limited by a high ASD of 1.45.
Recent methods such as PVT-CASCADE, PVT-GCASCADE-B2, and PVT-EMCAD-B2 demonstrate competitive performance, with PVT-CASCADE achieving strong results in Dice and IoU for all organs and the lowest ASD for the Aorta. PVT-EMCAD-B2 further improves upon these results, particularly in Heart and Trachea segmentation, achieving the highest IoU and competitive ASD values. However, the proposed EDLDNet method surpasses all other approaches, achieving the highest Dice and IoU scores for all organs and the lowest ASD values for the Heart, Trachea, and Aorta. For challenging regions like the Esophagus, it records significant improvements despite inherent segmentation difficulties due to its variable shape.
@@ 4.7 Results on LCTSC Dataset
The results in Table 4 compare segmentation performance across eight methods for the Esophagus, Spine, Heart, and Left and Right Lungs (Lung(L) and Lung(R)), evaluated using Dice score, ASD, and IoU. For the Esophagus, our method (EDLDNet) achieves the highest Dice (0.74) and IoU (0.87), significantly outperforming others, while AttUNet records the lowest ASD (1.09) but with lower Dice (0.66) and IoU (0.70), indicating weaker region segmentation. Spine segmentation shows consistent performance across methods, with Dice values of 0.88-0.89 and IoU of 0.91-0.92. AttUNet, UNet++, UDBANet, and our method achieve the lowest ASD (0.66), suggesting that the Spine is relatively easier to segment.
Recent methods such as PVT-CASCADE, PVT-GCASCADE-B2, and PVT-EMCAD-B2 demonstrate competitive performance across all organs. PVT-CASCADE achieves strong results in Heart segmentation, with a Dice score of 0.93 and the lowest ASD of 1.11, while PVT-EMCAD-B2 matches these results and further improves ASD for the Lungs. PVT-GCASCADE-B2 also performs well, achieving the lowest ASD for Lung(R) (0.64). However, our method consistently outperforms these approaches, achieving the highest Dice (0.94) and IoU (0.89) for the Heart, along with the lowest ASD (1.11). For the Lungs, our method demonstrates exceptional performance, with the highest Dice (0.97) and IoU (0.93) for both lungs and competitive ASD values (0.59 for Lung(L) and 0.64 for Lung(R)). While UDBANet matches our Dice scores for the lungs, its slightly higher ASD values reflect less precise boundary alignment. Other methods, such as AttUNet and UNet++, show competitive but less precise results, and R2UNet consistently underperforms. Overall, our method achieves state-of-the-art performance across all metrics and organs, demonstrating its robustness and precision in medical image segmentation.
@@ 5 Conclusion
CRediT authorship contribution statement
Riad Hassan: Conceptualization, Methodology, Investigation, Software, Formal analysis, Writing - Original draft. M. Rubaiyat Hossain Mondal: Conceptualization, Project Supervision, Writing review and editing. Sheikh Iqbal Ahamed: Project administration, Writing review and editing. Fahad Mostafa: Formal analysis, Writing review and editing. Md Mostafijur Rahman: Conceptualization, Software, Methodology, Validation, Writing review and editing.
Ethics Statement
This research did not involve any primary data collection or interaction with human or animal subjects. Consequently, formal ethical approval was not applicable. We affirm that this work has been conducted with rigorous adherence to academic standards and principles of ethical scholarly practice.

### v2

@@ 4.6 Result on SegThor Dataset
The comparison in Table 3 evaluates segmentation methods for four anatomical structures—Esophagus, Heart, Trachea, and Aorta—using metrics like Dice similarity, ASD, and mIoU. AttUNet performs well in ASD for the Esophagus but falls behind in Dice and mIoU, which indicates challenges with boundary segmentation accuracy. Similarly, R2UNet struggles across all metrics, especially for the Esophagus and Heart, reflecting its limitations in capturing complex anatomical details. Advanced methods like UNet++ and UDBANet show notable improvements, with UNet++ excelling in Heart and Trachea segmentation (Dice of 0.95 and 0.89) and UDBANet performing strongly for the Heart and Aorta, though its precision for the Esophagus is limited by a high ASD of 1.45.
Recent methods such as PVT-CASCADE, PVT-GCASCADE-B2, and PVT-EMCAD-B2 demonstrate competitive performance, with PVT-CASCADE achieving strong results in Dice and mIoU for all organs and the lowest ASD for the Aorta. PVT-EMCAD-B2 further improves upon these results, particularly in Heart and Trachea segmentation, achieving the highest mIoU and competitive ASD values. However, the proposed EDLDNet method surpasses all other approaches, achieving the highest Dice and mIoU scores for all organs and the lowest ASD values for the Heart, Trachea, and Aorta. For challenging regions like the Esophagus, it records significant improvements despite inherent segmentation difficulties due to its variable shape.
@@ 4.7 Results on LCTSC Dataset
The results in Table 4 compare segmentation performance across eight methods for the Esophagus, Spine, Heart, and Left and Right Lungs (Lung(L) and Lung(R)), evaluated using Dice score, ASD, and mIoU. For the Esophagus, the proposed method (EDLDNet) achieves the highest Dice (0.74) and mIoU (0.87), significantly outperforming others, while AttUNet records the lowest ASD (1.09) but with lower Dice (0.66) and mIoU (0.70), indicating weaker region segmentation. Spine segmentation shows consistent performance across methods, with Dice values of 0.88-0.89 and mIoU of 0.91-0.92. AttUNet, UNet++, UDBANet, and the proposed method achieve the lowest ASD (0.66), suggesting that the Spine is relatively easier to segment.
Recent methods such as PVT-CASCADE, PVT-GCASCADE-B2, and PVT-EMCAD-B2 demonstrate competitive performance across all organs. PVT-CASCADE achieves strong results in Heart segmentation, with a Dice score of 0.93 and the lowest ASD of 1.11, while PVT-EMCAD-B2 matches these results and further improves ASD for the Lungs. PVT-GCASCADE-B2 also performs well, achieving the lowest ASD for Lung(R) (0.64). However, the proposed method consistently outperforms these approaches, achieving the highest Dice (0.94) and mIoU (0.89) for the Heart, along with the lowest ASD (1.11). For the Lungs, the proposed method demonstrates exceptional performance, with the highest Dice (0.97) and mIoU (0.93) for both lungs and competitive ASD values (0.59 for Lung(L) and 0.64 for Lung(R)). While UDBANet matches our Dice scores for the lungs, its slightly higher ASD values reflect less precise boundary alignment. Other methods, such as AttUNet and UNet++, show competitive but less precise results, and R2UNet consistently underperforms. Overall, the proposed method achieves state-of-the-art performance across all metrics and organs, demonstrating its robustness and precision in medical image segmentation.
@@ 5 Conclusion
Acknowledgment
This work was partially supported by the Innovation Fund from ICT Division, Bangladesh under Reference No. 56.00.0000.053.33.0002.24.418.